# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
$\mathcal{A}_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{x_{3}}{x_{1}+\rho_{1}}<\tan\theta<\frac{x_{3}}{x_{1}-\rho_{1}},\,|\tan\phi|<\frac{h_{1}}{x_{3}},0\leq\phi\leq\pi\right\},$ (5.22)
$\mathcal{A}_{2}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{L-x_{3}}{d+\rho_{2}}<\tan\theta_{1}<\frac{L-x_{3}}{d-\rho_{2}},\,|\tan\phi_{2}|<\frac{h_{2}}{L-x_{3}},\pi\leq\phi_{2}\leq 2\pi\right\},$
$\mathcal{A}_{3}$ $=\mathbb{R}^{3}\setminus\left(\mathcal{A}_{1}\cup\mathcal{A}_{2}\right),$
where
$d$ $=\text{distance between}\,(x_{1},0,L)\,\text{and}\,a_{2},$
$\hat{r}_{1}$ $\coloneq x_{1}-x_{3}\cot\theta,\quad\hat{r}_{2}\coloneq d-(L-x_{3})\cot\theta,$
$h_{1}$ $\coloneq\sqrt{\rho^{2}_{1}-\hat{r}^{2}},\quad h_{2}\coloneq\sqrt{\rho^{2}_{2}-\hat{r}_{2}^{2}}.$
and $v=(v_{1},v_{2},v_{3})=(\rho\cos\theta,\rho\cos\phi\sin\theta,\rho\sin\phi\sin\theta)$ respect to standard basis for $\mathcal{A}_{1}$ but $v=(v_{1},v_{2},v_{3})=(\rho\cos\theta_{1},\rho\cos\phi_{1}\sin\theta_{1},\rho\sin\phi_{1}\sin\theta_{1})$ respect to the basis as
$e_{1}=\frac{x-a_{2}}{|x-a_{2}|},\quad\{e_{1},e_{2},e_{3}\}\,\text{is an orthogonal basis of }\mathbb{R}^{3}$
for $\mathcal{A}_{2}$. Also,
$\begin{aligned} \mathcal{A}_{1}&=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,0\leq\phi\leq\pi\right\}\quad&&\text{on }x_{3}=0,\\ \mathcal{A}_{2}&=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\pi\leq\phi\leq 2\pi\right\}\quad&&\text{on }x_{3}=L.\end{aligned}$
Note that $\mathcal{A}_{i}\subset A_{i}$ for any $i=1,2$. Define
$\partial\mathcal{D}$ $\coloneq\partial\left(\Omega\cap\{0\leq x_{3}\leq L\}\right)\setminus\Big\{\big(B(a_{1},2\rho_{1})\cap\{x_{3}=0\}\big)\cup\big(B(a_{2},2\rho_{2})\cap\{x_{3}=L\}\big)\Big\},$
$\partial\mathcal{A}_{3}$ $\coloneq\{(x,v)\in\partial\mathcal{D}\times\mathbb{R}^{3}|\,n(x)\cdot v<0\}.$
Notice that $\partial\mathcal{D}$ is not $C^{2}$, so the definition of $n(x)$ is taken from the original charts of $\partial\Omega$.
Now, since $\partial\mathcal{A}_{3}$ is a compact set in $\mathbb{R}^{3}$ and $\partial\Omega_{2}\subset\partial\mathcal{A}_{3}$. Similar to the proof of lemma 5.1, we can show that $\Omega$ is completely unsealed set with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$. ∎
The proof of Theorem 1.3 is similar to the proof of Theorem 1.2 because, after linearizing (1.22) by equation (5.2), (1.22) become
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=\frac{2}{\kappa}Q(m_{T_{q}},f)+\frac{1}{\kappa}Q(f,f)+\frac{1}{\kappa}Q(m_{T_{q}},m_{T_{q}}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,a},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,b},\\ f&=\frac{1}{2\pi T^{2}(x)}e^{-\frac{v^{2}}{2T(x)}}\int_{v\cdot n(x)>0}f(x,v)v\cdot n(x)dv,&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (5.23)
Then, equation (5.10) becomes
$\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv$ $=\int_{\mathbb{R}^{3}}v_{3}f(x,v)dv$ (5.24)
@@
Chapter 6. Appendix A. The sufficient conditions for unsealed sets
Lemma 6.1.
Proof.
Proposition 6.1.
  1. 1.
  2. 2.
  3. 3.
Then, $\Omega$ is unsealed with respect to $\partial\Omega_{1}$.
Proof.
For convenience, we use same symbol in Chapter 4
If $x\in\partial C_{2}$, then
Otherwise,
$=\displaystyle\int_{v\cdot n(x,v)>0}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(q(x,v))m(v)v\cdot n(x)dv$
$\leq\zeta.$
So
$\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{1})d\sigma_{0,1}\leq\zeta\quad\forall x\in\partial\Omega\setminus\partial\Omega_{1}.$ (6.1)
Chapter 7. Appendix B The equivalent definition of solution
Revision In this chapter, we discuss the definition of a solution of (1.5), (1.9) and valid that (5.9) exists when Theorem 1.1 holds. Also, we discuss the relation between the mild solution and $v\cdot\nabla f$.
Definition 7.1.
$F$ is an mild solution of (1.5) if
$+\frac{1}{\kappa}\int_{0}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
Let
We have another type of definition to solution of (1.5). That is,
Definition 7.2.
$F$ is an exponential solution of (1.5) if
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$. Here,
After linearizing (1.5), we obtain (1.9). Similar (1.5), we can define the mild solution as follow.
Definition 7.3.
$f$ is an mild solution of (1.9) if
$f$ $=\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}(q(x,v),v)$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
Also, we can define the exponential solution
Definition 7.4.
$f$ is an exponential solution of (1.9) if
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
It is obvious that the existence of a mild solution to (1.5) is equivalent to the existence of a mild solution to (1.9). Since the proof of the equivalence between Definition 7.1 and Definition 7.2 is similar to that between Definition 7.3 and Definition7.4, we only present the proof for the first case. The statement is as follows.
Theorem 7.1.
Suppose that $\nu_{F}(x-sv,v)$ and $Q_{+}(F,F)(x-sv,v)$ belong to $L^{1}_{loc}(-\tau_{-}(x,-v),\tau_{-}(x,v))$ for almost all $(x,v)\in\Omega\times\mathbb{R}^{3}$ and $F\in L^{1}(\Omega\times\mathbb{R}^{3})$. Then, the definition of 7.1 is equivalent to the definition of 7.2.
Proof.
Let
and
$G(s)=g(x,v)+\frac{1}{\kappa}\int_{s}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt,$
and
$\left\{\begin{aligned} \frac{d}{ds}G(s)&=-\frac{1}{\kappa}Q(F,F)(x-sv)\\ &=-\frac{1}{\kappa}Q_{+}(F,F)(x-sv)+\frac{1}{\kappa}\nu_{F}(x-sv)F(x-sv)\\ &=-\frac{1}{\kappa}Q_{+}(F,F)(x-sv)+\frac{1}{\kappa}\nu_{F}(x-sv)G(s),&&\text{ in }s\in[-\tau_{-}(x,-v),\tau_{-}(x,v)].\\ G(0)&=F(x,v).\end{aligned}\right$. (7.1)
Note that the solution of (7.1) is unique and $G$ has two representations. One of them is
$G(s)=g(x,v)+\frac{1}{\kappa}\int_{s}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt.$
The other is
By uniqueness, the proof is complete. ∎
Now, we consider the following boundary value problem.
[…]
We pay attention on the following question
The mild solution of (7.2) is
$F(x,v)=F_{in}(q(x,v),v)+\int_{0}^{\tau_{-}(x,v)}G(x-sv,v)ds.$
Theorem 7.2.
If $F$ is a mild solution of (7.2) and $F,G\in L^{p}(\Omega\times\mathbb{R}^{3})$ for some $p\geq 1$, then
for any $\varphi\in C^{\infty}_{c}(\Omega\times\mathbb{R}^{3})$.
Proof.
Notice that
$=\int_{0}^{h}G(x+sv,v)ds$
Moreover, we have
Hence the proof is complete. ∎
Chapter 8. The $H^{2}$ estimate of Laplace equation with Robin boundary condition
Revision In this chapter, we prove the inequality (2.12). That is, we discuss about the following elliptic equation
$\left\{\begin{aligned} -\Delta U&=F,&&\text{in }\Omega,\\ \hat{\gamma}(x)U+\partial_{\nu}U&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.1)
Here $\hat{\gamma}(x)$ is a non-negative $C^{2}$ function and $\hat{\gamma}(x)=0$ on $\partial\Omega_{2}$ but $\hat{\gamma}>0$ on a positive measure subset of $\partial\Omega$. Before we start our argument, we recall several classical trace theorem and elliptic estimate.
@@
    $\int_{\Omega}\Delta u\,dx=\int_{\partial\Omega}\partial_{\nu}u\,dS.$
@@
    $\int_{\Omega}\nabla v\cdot\nabla u\,dx=-\int_{\Omega}u\Delta v\,dx+\int_{\partial\Omega}u\,\partial_{\nu}v\,dS.$
@@
    $\int_{\Omega}\bigl(u\,\Delta v-v\,\Delta u\bigr)\,dx=\int_{\partial\Omega}\left(u\,\partial_{\nu}v-v\,\partial_{\nu}u\right)\,dS.$
@@
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ with $C^{2}$ boundary. Assume that $F\in L^{2}(\Omega)$. Suppose that $U\in H^{1}(\Omega)$ is a weak solution of (8.1). Then
@@
$\int_{\Omega}Fdx=\int_{\partial\Omega}\hat{\gamma}(x)UdS.$ (8.2)
@@
$\left\{\begin{aligned} -\Delta U&=F,&&\text{in }\Omega,\\ \partial_{n}U&=g\coloneq-\hat{\gamma}(x)U,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.3)
After applying theorem 8.3, there exists $G\in W^{2}(\Omega)$ such that $\partial_{\nu}G=-\hat{\gamma}(x)U$. Then, UGU-G satisfies
$\left\{\begin{aligned} -\Delta(U-G)&=F+\Delta G,&&\text{in }\Omega,\\ \partial_{n}(U-G)&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.4)
By (8.2) and Green’s identity, we have
$\int_{\Omega}F+\Delta Gdx=\int_{\Omega}Fdx+\int_{\Omega}\Delta Gdx=\int_{\Omega}Fdx+\int_{\partial\Omega}\partial_{\nu}GdS=\int_{\Omega}Fdx-\int_{\partial\Omega}\hat{\gamma}(x)UdS=0.$
@@
$\leq C\left(\|F\|_{L^{2}(\Omega)}+\|\hat{\gamma}(x)U\|_{H^{1/2}(\partial\Omega)}\right)$
@@
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ with $C^{2}$ boundary. Assume that $F\in L^{2}(\Omega)$. Then, (8.1) admits a unique weak solution $U$. Furthermore,
$\|U\|_{H^{1}(\Omega)}\leq C\|F\|_{L^{2}(\Omega)},$ (8.5)
@@
Note that the weak solution of (8.1) means
$\int_{\Omega}\nabla U\cdot\nabla v\,dx+\int_{\partial\Omega}\hat{\gamma}Uv\,dS=\int_{\Omega}Fv\,dx.$ (8.6)
@@
$a(u,v)\coloneq\int_{\Omega}\nabla u\cdot\nabla v\,dx+\int_{\partial\Omega}\hat{\gamma}uv\,dS.$
Since $\hat{\gamma}\in L^{\infty}(\partial\Omega)$ and $u,v\in H^{1}(\Omega)$, we have
@@
$\|u_{n}\|^{2}_{H^{1}(\Omega)}>n\left(\int_{\Omega}|\nabla u_{n}|^{2}\,dx+\int_{\partial\Omega}\hat{\gamma}u_{n}^{2}\,dS\right).$
Without loss of generality, we assume that $\|u_{n}\|_{L^{2}(\Omega)}=1$ for any $n\in\mathbb{N}$. We can discover that the weak limit of $\{u_{n}\}$ exists by above inequality. Let $u=w-\lim_{n\to\infty}u$. Observe that
$0=\lim_{n\to\infty}\frac{1}{n}\|\nabla u\|_{L^{2}(\Omega)}\geq\lim_{n\to\infty}\|\nabla u_{n}\|_{L^{2}(\Omega)}\|\nabla u\|_{L^{2}(\Omega)}\geq\lim_{n\to\infty}\int_{\Omega}\nabla u_{n}\cdot\nabla u\,dx=\int_{\Omega}|\nabla u|^{2}\,dx.$
It implies that there exists a subsequence of $u_{n}$ which strongly converge to some $u$ in $H^{1}(\Omega)$. Moreover, we deduce that $u=0$ and obtain a contradiction since $\|u_{n}\|_{L^{2}(\Omega)}=1$. By Lax-Milgram theorem, the proof of the existence is complete. Assume that (8.5) does not hold. There exists a sequence $\{U_{n}\}$ such that
$\|U_{n}\|_{H^{1}(\Omega)}\geq n\|F_{n}\|_{L^{2}(\Omega)}.$
@@
$\int_{\Omega}|\nabla U_{n}|^{2}\,dx+\int_{\partial\Omega}\hat{\gamma}U_{n}^{2}\,dS=\int_{\Omega}F_{n}U_{n}\,dx\leq\|F_{n}\|_{L^{2}(\Omega)}\|U_{n}\|_{L^{2}(\Omega)}.$
Without loss of generality, we still assume that $\|U_{n}\|_{L^{2}(\Omega)}=1$ for any $n\in\mathbb{N}$. After a similar argument, we can immediately conclude that $\{U_{n}\}$ is bounded in $H^{1}(\Omega)$ and there exists a subsequence of $U_{n}$ which strongly converge to 0. It is a contradiction. The uniqueness immediately comes from (8.6) and the similar argument. ∎
After applying the above estimate, we can prove the proposition 2.6 here.
@@
The Green’s identity provides the following weak version of (2.8):
$\int_{\partial\Omega\times\mathbb{R}^{3}}\psi f\,dSdv-\iint_{\Omega\times\mathbb{R}^{3}}v\cdot\nabla_{x}\psi\,fdxdv=-\iint_{\Omega\times\mathbb{R}^{3}}\psi\,\frac{1}{\kappa}L(\mathrm{I}-\mathbf{P})fdxdv+\iint_{\Omega\times\mathbb{R}^{3}}\psi\,\phi_{1}dxdv.$ (8.7)
We recall that
$P_{L}(f)=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}.$
The key of the proof is to choose suitable $H^{1}$ test functions $\psi$ to estimate $\langle f,\Phi_{i}\rangle_{v}$. We choose the test function
$\psi=(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0},$ (8.8)
$\left\{\begin{aligned} -\Delta\phi_{\langle f,\Phi_{0}\rangle_{v}}&=\langle f,\Phi_{0}\rangle_{v},&&\text{in }\Omega,\\ \hat{\gamma}(x)\phi_{\langle f,\Phi_{0}\rangle_{v}}+\partial_{n}\phi_{\langle f,\Phi_{0}\rangle_{v}}&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (8.9)
Here $\hat{\gamma}(x)$ is a non-negative $C^{2}$ function and $\hat{\gamma}(x)=0$ on $\partial\Omega_{2}$ ( but $\gamma>0$ on a positive measure subset of $\partial\Omega$ ). We bound the right hand side of (2.8) by
We choose $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ such that
$\int_{\mathbb{R}^{3}}\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)\left(\frac{|v|^{2}}{2}-\frac{3}{2}\right)(v_{i})^{2}\,\mu(v)\,dv=0.$ (8.10)
Moreover, we decompose that
$\begin{aligned} f&=\mathbf{1}_{\partial\Omega^{+}_{1}}f+\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}+\mathbf{1}_{\partial\Omega_{2}}Pf+\mathbf{1}_{\partial\Omega^{+}_{2}}(I-P)f+\mathbf{1}_{\partial\Omega^{-}_{2}}\phi_{2},&&\qquad\text{on }\partial\Omega\times\mathbb{R}^{3},\\ f&=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}+(\mathrm{I}-\mathbf{P}_{L})f,&&\qquad\text{on }\Omega\times\mathbb{R}^{3}.\end{aligned}$ (8.11)
By (8.11), since the $\langle f,\Phi_{4}\rangle_{v}$ contribution vanishes in (2.8) due to our choice of $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ and the $\langle f,\Phi_{1}\rangle_{v}$, $\langle f,\Phi_{2}\rangle_{v}$, and $\langle f,\Phi_{3}\rangle_{v}$ contribution vanish in (2.8) due to the oddness, the left-hand side of (2.8) takes the form of
$\int_{\partial\Omega\times\mathbb{R}^{3}}n\cdot v\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)v\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0}\bigl[\mathbf{1}_{\partial\Omega^{+}_{1}}f+\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}+\mathbf{1}_{\partial\Omega_{2}}Pf+\mathbf{1}_{\partial\Omega^{+}_{2}}(I-P)f+\mathbf{1}_{\partial\Omega^{-}_{2}}\phi_{2}\bigr]$ (8.12)
$-\sum_{i,k=1}^{d}\iint_{\Omega\times\mathbb{R}^{3}}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{i}v_{k}\,\partial_{ik}\phi_{\langle f,\Phi_{0}\rangle_{v}}\,\langle f,\Phi_{0}\rangle_{v}\Phi^{2}_{0}dxdv$ (8.13)
$-\sum_{i,k=1}^{d}\iint_{\Omega\times\mathbb{R}^{3}}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{i}v_{k}\,\partial_{ik}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0}\,(\mathrm{I}-\mathbf{P}_{L})f.$ (8.14)
$v=(v\cdot n)n+v-(v\cdot n)n=v_{n}n+v_{\perp}.$
The contribution of $Pf=z(x)m^{\frac{1}{2}}=z(x)\Phi_{0}$ is
The crucial choice of $\psi$ makes the first term vanish due to the Neumann boundary condition, while the second term also vanishes due to the oddness of $v_{n}v_{\perp}$. Therefore, (8.12) and (8.14) are bounded by

### v2

@@ 4.1. The existence of solution to (80)
$W_{2}(x)\coloneq\int_{u\cdot n(x)>0}1_{\partial(\Omega\cap\{0\leq x_{3}\leq L\})\setminus A_{2}}(q(x,u))m_{T(q(x,u))}u\cdot n(x)du=1-\int_{u\cdot n(x)>0}1_{A_{2}}(u)m_{T(q(x,u))}u\cdot n(x)du.$
Since $\Omega\cap\{0\leq x_{3}\leq L\}$ is a convex set and $\partial\Omega\cap\{0\leq x_{3}\leq L\}$ is a $C^{1}$ boundary, $W_{1}$ is continuous on $\partial\Omega\cap[\frac{L}{2},L]$ and $W_{2}$ is continuous on $\partial\Omega\cap[0,\frac{L}{2}]$. Moreover, $W_{1}$ and $W_{2}$ are continuous on $\overline{\Omega\cap\{x_{3}=L\}\setminus\Pi}$ and $\overline{\Omega\cap\{x_{3}=0\}\setminus\Pi}$ respectively. Hence, we have
$\sup_{x\in\Xi\setminus\Pi}\int_{u\cdot n(x)>0}1_{\Xi\setminus\Pi}(q(x,u))m_{T(q(x,u))}u\cdot n(x)du$
$\leq\max\left\{\max_{x\in\partial\Omega\cap[\frac{L}{2},L]}W_{1}(x),\max_{x\in\overline{\Omega\cap\{x_{3}=L\}\setminus\Pi}}W_{1}(x),\max_{x\in\partial\Omega\cap[0,\frac{L}{2}]}W_{2}(x),\max_{x\in\overline{\Omega\cap\{x_{3}=0\}\setminus\Pi}}W_{2}(x)\right\}$
$\coloneq\zeta.$
Similar to the argument of lemma 5.1, we deduce that $\zeta$ is less than one. Thus, we conclude that $\Omega$ is completely unsealed set with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$. ∎
The proof of Theorem 1.3 is similar to the proof of Theorem 1.2 because, after rewriting (24) by equation (136), (24) becomes
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=\frac{2}{\kappa}Q(m_{T_{q}},f)+\frac{1}{\kappa}Q(f,f)+\frac{1}{\kappa}Q(m_{T_{q}},m_{T_{q}}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,a},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,b},\\ f&=\frac{1}{2\pi T^{2}(x)}e^{-\frac{v^{2}}{2T(x)}}\int_{v\cdot n(x)>0}f(x,v)v\cdot n(x)dv,&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (159)
Then, equation (144) becomes
$\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv$ $=\int_{\mathbb{R}^{3}}v_{3}f(x,v)dv$ (160)
@@
Acknowledgments
Prof. Tai-Ping Liu kindly discussed part of this problem with us. The authors would like to thank him for his interest and insightful suggestions. Also, the author would like to thank researcher Jhe-Kuan Su for giving fruitful comments. The research of the authors is supported in part by NSTC Grant 112-2115-M-002-009-MY3.
Section 6. Appendix A: The $H^{2}$ estimate of Laplace equation with Robin boundary condition
In this section, we prove the inequality (39). That is, we discuss about the following elliptic equation
$\left\{\begin{aligned} -\Delta U&=F,&&\text{in }\Omega,\\ \chi(x)U+\partial_{\nu}U&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (161)
Here $\chi(x)$ is a non-negative $C^{2}$ function and $\chi(x)=0$ on $\partial\Omega_{2}$. We require that $\chi>0$ on a subset of $\partial\Omega$ of positive measure. Before we start our argument, we recall several classical trace theorems and elliptic estimate. Here $\gamma$ is the trace operator.
@@ Acknowledgments
    $\int_{\Omega}\Delta u\,dx=\int_{\partial\Omega}\gamma\partial_{\nu}u\,dS.$
@@
    $\int_{\Omega}\nabla v\cdot\nabla u\,dx=-\int_{\Omega}u\Delta v\,dx+\int_{\partial\Omega}u\,\gamma\partial_{\nu}v\,dS.$
@@
    $\int_{\Omega}\bigl(u\,\Delta v-v\,\Delta u\bigr)\,dx=\int_{\partial\Omega}\left(u\,\gamma\partial_{\nu}v-v\,\gamma\partial_{\nu}u\right)\,dS.$
@@
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ with $C^{2}$ boundary. Assume that $F\in L^{2}(\Omega)$. Suppose that $U\in H^{1}(\Omega)$ is a weak solution of (161). Then
@@
$\int_{\Omega}Fdx=\int_{\partial\Omega}\chi(x)UdS.$ (162)
@@
$\left\{\begin{aligned} -\Delta U&=F,&&\text{in }\Omega,\\ \partial_{n}U&=g\coloneq-\chi(x)U,&&\text{on }\partial\Omega.\end{aligned}\right$. (163)
After applying theorem 6.3, there exists $G\in W^{2}(\Omega)$ such that $\partial_{\nu}G=-\chi(x)U$. Then, UGU-G satisfies
$\left\{\begin{aligned} -\Delta(U-G)&=F+\Delta G,&&\text{in }\Omega,\\ \partial_{n}(U-G)&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (164)
By (162) and Green’s identity, we have
$\int_{\Omega}F+\Delta Gdx=\int_{\Omega}Fdx+\int_{\Omega}\Delta Gdx=\int_{\Omega}Fdx+\int_{\partial\Omega}\partial_{\nu}GdS=\int_{\Omega}Fdx-\int_{\partial\Omega}\chi(x)UdS=0.$
@@
$\leq C\left(\|F\|_{L^{2}(\Omega)}+\|\chi(x)U\|_{H^{1/2}(\partial\Omega)}\right)$
@@
Let $\Omega$ be a bounded connected open subset of $\mathbb{R}^{n}$ with $C^{2}$ boundary. Assume that $F\in L^{2}(\Omega)$. Then, (161) admits a unique weak solution $U$. Furthermore,
$\|U\|_{H^{1}(\Omega)}\leq C\|F\|_{L^{2}(\Omega)},$ (165)
@@
Note that the weak solution of (161) means
$\int_{\Omega}\nabla U\cdot\nabla v\,dx+\int_{\partial\Omega}\chi Uv\,dS=\int_{\Omega}Fv\,dx.$ (166)
@@
$a(u,v)\coloneq\int_{\Omega}\nabla u\cdot\nabla v\,dx+\int_{\partial\Omega}\chi uv\,dS.$
Since $\chi\in L^{\infty}(\partial\Omega)$ and $u,v\in H^{1}(\Omega)$, we have
@@
$\|u_{n}\|^{2}_{H^{1}(\Omega)}>n\left(\int_{\Omega}|\nabla u_{n}|^{2}\,dx+\int_{\partial\Omega}\chi u_{n}^{2}\,dS\right).$ (167)
Without loss of generality, we assume that $\|u_{n}\|_{L^{2}(\Omega)}=1$ for any $n\in\mathbb{N}$. By (167) and $\|u_{n}\|_{L^{2}(\Omega)}=1$, we discover that $\{u_{n}\}$ is a bounded sequence in $H^{1}(\Omega)$. Hence $\{u_{n}\}$ is weakly compact in $H^{1}(\Omega)$ and is compact in $L^{2}(\Omega)$. Let $\{u_{n_{j}}\}$ be a convergent subsequence of $\{u_{n}\}$ and $u$ be its limit. That is,
$\begin{aligned} &u_{n_{j}}\rightharpoonup u&&\quad\text{ in }H^{1}(\Omega),\\ &u_{n_{j}}\to u&&\quad\text{ in }L^{2}(\Omega).\end{aligned}$ (168)
We observe that
$0=\lim_{n_{j}\to\infty}\frac{1}{n_{j}}\|\nabla u\|_{L^{2}(\Omega)}\geq\lim_{n_{j}\to\infty}\|\nabla u_{n_{j}}\|_{L^{2}(\Omega)}\|\nabla u\|_{L^{2}(\Omega)}\geq\lim_{n_{j}\to\infty}\int_{\Omega}\nabla u_{n_{j}}\cdot\nabla u\,dx=\int_{\Omega}|\nabla u|^{2}\,dx.$
It implies that $u_{n_{j}}$ converges strongly to $u$ in $H^{1}(\Omega)$; hence,
$\lim_{n_{j}\to\infty}\int_{\partial\Omega}\chi u_{n}^{2}\,dS=\int_{\partial\Omega}\chi u^{2}\,dS=0.$
Thus, we deduce that $u=0$, which yields a contradiction since $\|u_{n}\|_{L^{2}(\Omega)}=1$. By Lax-Milgram theorem, the proof of the existence is complete. Assume that (165) does not hold, there exists a sequence $\{U_{n}\}$ such that
$\|U_{n}\|_{H^{1}(\Omega)}\geq n\|F_{n}\|_{L^{2}(\Omega)}.$ (169)
@@
$\int_{\Omega}|\nabla U_{n}|^{2}\,dx+\int_{\partial\Omega}\chi U_{n}^{2}\,dS=\int_{\Omega}F_{n}U_{n}\,dx\leq\|F_{n}\|_{L^{2}(\Omega)}\|U_{n}\|_{L^{2}(\Omega)}.$ (170)
Without loss of generality, we assume that $\|U_{n}\|_{L^{2}(\Omega)}=1$ for any $n\in\mathbb{N}$. By (169) and (170), we can immediately conclude that $\{U_{n}\}$ is bounded in $H^{1}(\Omega)$ and there exists a subsequence of $U_{n}$ which strongly converge to 0 in $H^{1}(\Omega)$ by similar argument. It is a contradiction since $\|U_{n}\|_{L^{2}(\Omega)}=1$ for any $n\in\mathbb{N}$. The proof is complete. ∎
Section 7. Appendix B: The sufficient conditions for unsealed sets
Lemma 7.1.
@@
Proposition 7.1.
  1. 1.
  2. 2.
  3. 3.
Then, $\Omega$ is unsealed with respect to $\partial\Omega_{1}$.
Proof.
For convenience, we use same symbol in section 4
If $x\in\partial C_{2}$, then
Otherwise,
$=\int_{v\cdot n(x,v)>0}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(q(x,v))m(v)v\cdot n(x)dv$
$\leq\zeta.$
So
$\displaystyle\int\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{2})d\sigma_{2,1}\textbf{1}_{\partial\Omega\setminus\partial\Omega_{1}}(x_{1})d\sigma_{0,1}\leq\zeta\quad\forall x\in\partial\Omega\setminus\partial\Omega_{1}.$ (171)
Section 8. Appendix C: The equivalent definition of solution
In this section, we discuss the definition of a solution of (7), (11) and valid that (143) exists when Theorem 1.1 holds. Also, we discuss the relation between the mild solution and $v\cdot\nabla f$.
Definition 8.1.
$F$ is an mild solution of (7) if
$+\frac{1}{\kappa}\int_{0}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
Let
We have another type of definition to solution of (7). That is,
Definition 8.2.
$F$ is an exponential solution of (7) if
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$. Here,
After linearizing (7), we obtain (11). Similar (7), we can define the mild solution as follow.
Definition 8.3.
$f$ is an mild solution of (11) if
$f$ $=\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}(q(x,v),v)$
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
Also, we can define the exponential solution
Definition 8.4.
$f$ is an exponential solution of (11) if
a.e. $(x,v)\in\Omega\times\mathbb{R}^{3}$.
It is obvious that the existence of a mild solution to (7) is equivalent to the existence of a mild solution to (11). Since the proof of the equivalence between Definition 8.1 and Definition 8.2 is similar to that between Definition 8.3 and Definition8.4, we only present the proof for the first case. The statement is as follows.
Theorem 8.1.
Suppose that $\nu_{F}(x-sv,v)$ and $Q_{+}(F,F)(x-sv,v)$ belong to $L^{1}_{loc}(-\tau_{-}(x,-v),\tau_{-}(x,v))$ for almost all $(x,v)\in\Omega\times\mathbb{R}^{3}$ and $F\in L^{1}(\Omega\times\mathbb{R}^{3})$. Then, the definition of 8.1 is equivalent to the definition of 8.2.
Proof.
Let